# Numerical analysis of heat transport in classical one-dimensional systems

### v1

@@ 7 Conclusion
Fluctuating hydrodynamics shows that three conservation laws (energy, linear momentum, and stretch) are a prerequisite for the emergence of anomalous heat conductivity in one dimension. Several publications have anyhow challenged this scenario, providing examples of a finite conductivity in models where a divergence should be instead observed. I have revisited some of them, showing that these previous conclusions are based on finite-size effects: (much) longer system-sizes are needed in order to see the divergence predicted by theory. The observation of a finite conductivity in the model of Ref. [20] is an exception, presumably due to an error in the estimate of the total flux.
Now, that several inconsistencies between predictions and observartions have been removed, a question still stands about the emergence, in some models, of a seemingly finite conductivity over a broad range of scales. The answer is straightforward while dealing with nearly integrable systems, but it is much less so when the model itsels is manifestly chaotic, as it happens for the soft point chain analysed in this paper. Moreover, not all non-binding potentials exhibit a nearly constant effective conductivity.
An additional complication comes from the amplitude of term gauging the strength of the diverging contribution; it is very small in the homogeneous soft point chain, while it diverges in the stochastic linear model when it approaches the integrable limit. How to justify such a wide range of different results?

### v2

@@ 6 Conclusion
Fluctuating hydrodynamics shows that three conservation laws (energy, linear momentum, and stretch) are a prerequisite for the emergence of anomalous heat conductivity in one dimension. Several publications have anyhow challenged this scenario, providing examples of finite conductivity in models where a divergence should be instead observed. I have revisited some of them, showing that in all cases heat conductivity diverges in the thermodynamic limit. In the case of the ding-a-ling type model studied in Ref. [19], the discrepancy is due to some kind of error presumably related to the definition of the local flux. More interesting is the case of non-binding potentials, where I find that the conductivity is affected by very strong finite-size effects, which mask the asymptotic behavior. This conclusion is supported by the observation that the effective conductivity can be well approximated as the sum of two distinct contributions: a kinetic diffusive term and the anomalous one. This assumption, previously proposed to explain the behavior of nearly integrable systems, proves, indeed, more general and is valid even in fully chaotic models,
As a result, a question opens up: how can it be that some modes can propagate over very long distances in the presence of an evident non-integrability? In a sense, this unexpected scenario is complentary to the anomalous (non ballistic) divergence observed in the symmetric gard-point chain, where the integrals of motions (the velocitons) diffuse rather than propagating ballistically. An explanation of when and how this scenario emerges in non-binding potential is a point that will be worth exploring.
Finally, while representing the conductivity as the sum of two terms, I have also found that the amplitude $\sigma$ of the hydrodynamic term plays a crucial role in determining the crossover length. In one of the non-binding potentials, it is very small, thus contributing to further increase the crossover length. Understanding which dynamical properties are responsible for the amplitude of $\sigma$ is another point that should be clarified.